y + y - 2y = e. y + y - 2y = €, y(0) = 0, /(0) =
Answer #1

0 0 Given ODE y+yl-2y = ex dy + de la zy zek By+ Dy-2y = ek (8248-2) y = 0 het f(D)= D?+ D-2 A.E. RS semo 1927 m-220 on=11-26 from @ @ . . 24 General solution of gltul-2y = e es Y = 9 (2) = 48 tee2 + z ea ② den Gpuen 4(0)=0 ve y=0 when an o = c +cat